OpenStudy (anonymous):

Evaluate the integral from 0 to 3 of the quotient of x and the square root of the quantity x squared plus 16, dx. a: 1 b: 2 c: 3 d: 4

OpenStudy (anonymous):

this? \(\huge \int_{0}^{3}\frac{x}{\sqrt{x^2+16}}\)

OpenStudy (anonymous):

if that is so, then the variable change \(x^2+16=u\) and \(2xdx=du\) should do the job

OpenStudy (anonymous):

so after the substitution you get: \(\huge\frac{1}{2}\int_{16}^{25}\frac{du}{\sqrt{u}}\)

OpenStudy (anonymous):

which integrated directly

OpenStudy (anonymous):

yeah with a dx to the right of it :)

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

@Nnesha @miss.happy @Michele_Laino

OpenStudy (michele_laino):

I think that the result of @myko is right

OpenStudy (anonymous):

the options are 1, 2, 3, or 4 :)

OpenStudy (michele_laino):

please use this identity: \[\int {\frac{{du}}{{\sqrt u }}} = 2\sqrt u \]

OpenStudy (anonymous):

what would i get as an answer? btw when myko asked: "this?" it was missing a dx next to it :)

OpenStudy (michele_laino):

Please I can not give you the answer directly since the Code of Conduct

OpenStudy (anonymous):

okay can u plz write me the steps then? :)

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

please you have to compute this: \[\left. {\sqrt u } \right|_{16}^{25}\]

OpenStudy (anonymous):

how do i do so? :)

OpenStudy (michele_laino):

since we can write: \[\frac{1}{2}\int {\frac{{du}}{{\sqrt u }}} = \sqrt u \]

OpenStudy (michele_laino):

hint: \[\left. {\sqrt u } \right|_{16}^{25} = \sqrt {25} - \sqrt {16} = ...?\]

OpenStudy (anonymous):

1 :) thank u !! i have 2 more if u dont mind?

OpenStudy (michele_laino):

ok!